OpenStudy (anonymous):

What is the exact value of the expression 11SqrT(8) + 6SqrT(12) − 5SqrT(2) ? Show your work.

OpenStudy (anonymous):

What can you tell me?

OpenStudy (anonymous):

That i am highly confused... Thats about it. I don't know anything about this problem.

OpenStudy (anonymous):

1) We cannot combine the terms unless the sqrts are all the same sqrt. Only sqrt 2 terms can be added to other sqrt 2 terms. We add the part outside the radical. 2) We should try to simplify each term individually to see if we can find some common ground. Can you simplify the terms individually?

OpenStudy (anonymous):

8=2^3 12=2^2(3) 2=2

OpenStudy (anonymous):

Think: \[\sqrt{8}=\sqrt{4 x 2} = 2\sqrt{2}\] \[\sqrt{12}=\sqrt{4 x 3} = 2\sqrt{3}\] Do you get me?

OpenStudy (anonymous):

Yes Sir.

OpenStudy (anonymous):

Now we have some sqrt(2) terms we can add, but don't forget to multiply the original outer numbers by the new outer numbers. What do we have now?

OpenStudy (anonymous):

\[4\sqrt{6}\]...?

OpenStudy (anonymous):

I get: \[22\sqrt{2} + 12\sqrt{3} - 5\sqrt{2}\] 22 - 5 = 17 so..... \[17\sqrt{2} + 12\sqrt{3}\] since I cannot combine the two different sqrts

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

Do you see where you went wrong? That is the most important thing. I want you to GET the answer more than me just giving to you.

OpenStudy (anonymous):

Yes Sir, i see where i went wrong with it. Thank You for the help.

OpenStudy (anonymous):

You are welcome.